Bronze which costs $9.10/kg is made by
combining copper which costs $8.90/kg with
tin which costs $9.50/kg. Find the number
of kg of copper and tin required to make
15.3 kg of bronze.

Answer by mananth(16949)

Copper x kg-------- $8.9
Tin y kg--------$9.5

1 x + 1 y = 15.3 .............1
8.9 x + 9.5 y = 139.23 .............2 (9.1*15.3)
Eliminate y
multiply (1)by -9.5
Multiply (2) by 1
-9.5 x -9.5 y = -145.35
8.9 x 9.5 y = 139.23
Add the two equations
-0.6 x = -6.12
/ -0.6
x = 10.20 kg
plug value of x in (1)
1 x + 1 y = 15.3
10.2 + 1 y = 15.3
1 y = 15.3 -10.2
1 y = 5.1
y = 5.10kg